# Many LLMs Are More Utilitarian Than One

### v1

@@ Abstract
Moral judgment is integral to large language model (LLM) alignment and social reasoning. As multi-agent systems gain prominence, it becomes crucial to understand how LLMs function collectively during collaboration, compared to individual agents. In human moral judgment, group deliberation leads to a utilitarian boost: a tendency to endorse norm violations that maximize benefits for the greatest number of people despite harms. We study whether a similar dynamic emerges in multi-agent LLM systems. We tested six models on well-established sets of moral dilemmas across two conditions: (1) Solo, where models reasoned independently, and (2) Group, where they engaged in multi-turn discussions in pairs or triads. In personal moral dilemmas, where agents must decide to directly harm one individual to maximize the utility for others, all models found moral violations to be more acceptable when part of a group than individually, similar to human experiments. Some models endorsed actions that maximized overall well-being, even if they benefited strangers over familiar individuals. Others became more willing to violate moral norms in groups. However, while human groups show a similar action bias, the mechanism for their utilitarian boost differs from LLMs. Whereas the human shift comes from heightened sensitivity to decision outcomes, LLM groups show either reduced norm sensitivity or enhanced impartiality. This suggests that while the surface behavior of LLM collectives mimics human group reasoning, the underlying drivers differ. We discuss the implications for AI alignment, multi-agent design, and artificial moral reasoning.
@@ 1 Introduction
Multi-agent systems (MAS) provide a key paradigm for analyzing decentralized decision-making and coordination among multiple autonomous entities in dynamic environments. By incorporating large language models (LLMs), these systems, referred to as LLM-MAS, feature agents with enhanced capabilities in reasoning, communication, and knowledge access [1]. This integration significantly expands the scope and complexity of tasks MAS can tackle, supporting more adaptive, flexible, and human-aligned agent behaviors [2]. However, integrating LLMs into MAS also introduces a new class of risks. When agents negotiate through iterative message passing or refine solutions jointly, even minor alignment discrepancies can cascade into major emergent distortions. Bias amplification [3], covert coordination protocols [4], knowledge drift [5], conflicting agreements [6], and collusion [7, 8] may emerge, all of which remain invisible when each LLM agent is evaluated in isolation. As multiple LLM-based assistants collaborate to tackle increasingly complex tasks, understanding how collective reasoning emerges from individual models is imperative for ensuring safety and trustworthiness [9].
@@
Figure 1: A schematic representing our experimental setup for LLM moral deliberation and reflection. A triad of of LLM agents engage in multi-round discussions about moral dilemmas and conclude with private reflections. This example illustrates how the group setting induces a utilitarian boost even when such actions involve moral norm violations.
@@
Recent work has begun to draw on insights from social psychology to investigate emergent distortions in group-level multi-agent LLM reasoning, demonstrating phenomena such as conformity [10, 11], belief congruence [12], and social irrationality [4]. Although not unique to LLM-MAS, these emergent patterns often show both similarities to and differences from well-studied collective behaviors in human groups [13]. Such collective distortions become especially important in moral domains where real-world consequences of norm violations can be disastrous, as in healthcare, education, and law. Aligning individual models will be insufficient to guarantee aligned group outcomes in such cases, and debiasing would need to take into account the collective element of the decision-making. But past research on LLM moral reasoning has overwhelmingly focused on comparing isolated agents with humans, as in the Moral Machine Experiment and the Moral Reasoning Benchmarks [14, 15, 16]. Thus, it provides little guidance for the emergent collective moral dynamics.
@@
This knowledge gap increases the risk of missing important deviations from normative guidelines, with strong possible real-world consequences. Moral decision-making is increasingly delegated to LLM collectives across high-stakes domains—for instance, coordinating specialized doctor and patient agents in simulated clinical settings [17], and decomposing complex legal tasks for improving legal reasoning and rule interpretation [18]. In these settings, covert consensus protocols could lead to collective LLM decisions that contradict those of reasonable humans, any normative frameworks, and even the output of any single model. This creates an urgent blind spot: without analyzing group-level moral dynamics in LLM-MAS, we cannot understand, predict, or prevent ethically problematic outcomes that escape single-agent safety evaluations. Suppose, for instance, that clinical LLM groups exhibit an emergent willingness to sacrifice individuals for “the greater good” in contrast with a physician’s Hippocratic oath. The resulting loss of trust in experts and decision-making institutions could be potentially catastrophic. Studying group dynamics in moral LLM decision-making also fills an important gap in computational social science, where the potential of LLMs for simulating human reasoning or determining whether human tendencies arise from language use is receiving greater attention [19]. Ignoring the deeply social nature of moral reasoning—where social deliberation, argumentation, and negotiation play a central role—also limits the insights afforded by simulation experiments.
This study applies controlled experiments and validated measurement tools from psychology to address the discussed gaps by systematically investigating how collective moral reasoning evolves in LLM-MAS. In doing so, it focuses on three central questions:
@@
    Are LLMs more likely to violate moral norms in the service of “the greater good” when deliberating in small groups? We operationalize this “utilitarian boost” by placing multiple copies of the same LLM in a structured dialogue, asking them to reach consensus on diverse sets of well-established moral scenarios.
@@
    How do multi-agent LLM judgments compare to human group judgments? Drawing on established human data on group deliberation studies, we examine both quantitative shifts (e.g., change in utilitarian endorsement rates) and the possible mechanisms at play.
@@
    How do any shifts or differences arise? We compare theories from moral psychology as explanations, seeking evidence for outcome-focused reasoning versus norm-violation rationalizations. We test multiple LLMs to assess mechanistic variability [^1] Mechanism refers to patterns in LLMs’ outputs across conditions, not architectures or cognitive processes.
As a preview, our experiments reveal a concerning phenomenon that has not been demonstrated in multi-agent LLM systems before: when identical agents engage in collective deliberation, they do not merely aggregate their individual moral judgments. Instead, they exhibit a consistent utilitarian boost—a systematic shift toward endorsing actions that maximize overall welfare, even when such actions involve sacrificing or harming a minority of humans. After six rounds of inter-agent exchange, groups reach a utilitarian decision that is more approving of moral norm violations than the corresponding single model. This utilitarian boost appears in six state-of-the-art LLMs and is therefore unlikely to reflect specific training datasets or paradigms. Reliability checks through human ratings make incoherent responses or misalignment between moral ratings and argumentation unlikely explanations for the change. While a similar utilitarian boost has been observed in human groups [20, 21, 22], we find the mechanisms for the LLM-MAS shift to vary across models and be markedly different from humans in most cases. Instead of reflecting the human groups’ focus on shared values, deliberative norms, and sensitivity to decision outcomes, the LLM-MAS shift may reflect increased impartiality for some models and a range of action preferences in others.
The utilitarian boost in collective LLM moral decision-making is particularly concerning given that LLMs are already deployed in ethically-sensitive domains, offering guidance in emotionally complex situations to resolve dilemmas involving fairness, harm, and moral responsibility [23, 24, 25], with some researchers urging their broader deployment by claiming that LLMs can outperform professional ethicists in moral judgment tasks [26]. To aid future investigations of what discursive elements facilitate this shift, we report in the supplementary materials an array of text-based analyses performed on LLM-generated arguments. Inspired by the literature on discursive predictors of utilitarianism and group influence in humans, these measures include harm-based argumentation, abstraction, emotionality, and conversational receptiveness. Overall, we find that these predictors of the utilitarian shift in humans do not explain the change in LLM collective judgments, highlighting the need for further investigations of how textual contexts affect LLMs’ downstream moral judgments.
@@ 2 Related Work
2.1 Emergent Biases in LLM Driven Multi Agent Systems
Insights from social psychology are increasingly being applied to analyze emergent reasoning biases in LLM-MAS. Liu et al. [4] introduce CogMir, a framework that leverages systematic hallucinations in LLMs to model human cognitive biases and assess prosocial irrationality in collective decision-making. Their findings reveal both parallels and divergences between human and LLM group behavior, particularly in how agents express prosocial tendencies under uncertainty. Weng et al. [10] investigate conformity among LLM agents, showing that models tend to align their responses with peers under social influence, mirroring human-like conformity effects. Borah et al. [12] examine belief congruence in LLM collectives and show that models are prone to belief alignment effects, which can amplify bias and distort reasoning in group contexts. While these studies highlight emergent group behaviors in LLM collectives, they largely overlook the moral domain. Human groups are known to exhibit distinct moral shifts, such as the utilitarian boost, during collective deliberation. Understanding whether similar effects arise in LLM-based systems is essential for assessing their alignment, especially in ethically sensitive applications.
@@ 2.2 Moral Reasoning in Human Groups
When humans act as a group, their collective moral reasoning cannot be fully explained in terms of the aggregate of the attitudes of individual group members. Empirical attempts have been made to show the direction and the mechanisms of this shift. When decisions are made collectively, people more readily accept breaches that increase the benefit of many [20, 21, 22]. One study identified heightened rationality in groups as the mechanism for the increased acceptance of moral breaches [20]. However, a different study found that the joint allocation of limited resources following a group discussion leads to less egalitarian distributions that benefit the least well-off [27]. Another study has proposed the reduction of stress and negative emotions as the potential underlying mechanism [21]. When social connections are induced between group members, pairs are more likely to jointly agree about sacrificing one to save many, providing yet another potential explanation for “the utilitarian boost” [28]. To adjudicate between these accounts, one recent study used computational modeling [29] to investigate whether groups’ greater utilitarianism stems from (1) stronger sensitivity to consequences, (2) weaker sensitivity to moral norms, or (3) weaker action aversion—and found that the utilitarian boost in human groups is driven solely by heightened sensitivity to consequences [22]. In other words, groups made more utilitarian choices not because they disregard norms or fear violating actions less, but because they are simply more sensitive to the consequences of their decisions . The current study uses the same modeling framework to examine whether the mechanisms of any utilitarian boosts are similar in LLMs.
@@ 3.1 Experimental Design
[…]
@@ 3.2 Stimuli and Measures
In line with human experiments, we use the Moral Utilitarian Dilemmas as our primary reference point to investigate the utilitarian boost in group or multi-agent settings [20, 21]. Drawn from the classic literature on moral judgments, the dilemmas assess how much agents prioritize collective outcomes over normative rules when faced with difficult ethical tradeoffs. The moral scenarios are compared with a non-moral baseline that accounts for generic response preferences. They vary along two key dimensions:
  • Personal vs. Impersonal [37]: personal dilemmas involve direct, emotionally charged harm (e.g. physically pushing one person to save five), whereas impersonal dilemmas involve indirect or mechanistic interventions (e.g. flipping a switch which will kill one to save many).
  • Action vs. Inaction [21]: action scenarios require a harmful intervention to achieve a “greater good”, while inaction scenarios describe situations where not doing anything enables harm while simultaneously generating benefits (e.g. withhold the truth from a close friend about their partner’s betrayal, preserving the relationship at the cost of induced harm).
We expected LLMs to be more sensitive to personal, active harm due their safety protocols. Once we established a utilitarian shift in this set, we used scenarios that orthogonally vary theoretically important attributes to disentangle whether it is driven by differential sensitivity to outcomes, a generic tendency to act in a moral context, or changed adherence to underlying moral obligations:
  • Oxford Utilitarianism Scale [38]: identifies Impartial Beneficence (IB) in the endorsement of utilitarian solutions. IB marks equal concern for all persons, contrasted with the willingness to deal direct harm (Instrumental Harm) as a potential cause for norm violation.
  • Factual Moral Framework [39]: comprises 32 real-world moral dilemmas, each varying by factual context (killing, other harm, or saving) and moral orientation (utilitarian vs. deontological). In our analysis we focus on three types: Killing–Utilitarian (sacrifice one to save many) and Other–Utilitarian (permit non-lethal harm for the greater good).
  • CNI Questionnaire [29]: this questionnaire is used to estimate three latent variables of utilitarian decision-making: 1) Consequence Sensitivity (C), or the weight the subject places on aggregate welfare, 2) Norm Sensitivity (N), or the adherence to moral rules, and 3) Inaction Preference (I), or the subject’s bias toward passive choices.
In the next section, we first examine the utilitarian shift across LLMs and group sizes, and later investigate possible mechanisms using these scales.

### v2

@@ Abstract
Moral judgment is integral to large language models’ (LLMs) social reasoning. As multi-agent systems gain prominence, it becomes crucial to understand how LLMs function when collaborating compared to operating as individual agents. In human moral judgment, group deliberation leads to a Utilitarian Boost: a tendency to endorse norm violations that inflict harm but maximize benefits for the greatest number of people. We study whether a similar dynamic emerges in multi-agent LLM systems. We test six models on well-established sets of moral dilemmas across two conditions: (1) Solo, where models reason independently, and (2) Group, where they engage in multi-turn discussions in pairs or triads. In personal dilemmas, where agents decide whether to directly harm an individual for the benefit of others, all models rated moral violations as more acceptable when part of a group, demonstrating a Utilitarian Boost similar to that observed in humans. However, the mechanism for the boost in LLMs differed: While humans in groups become more utilitarian due to heightened sensitivity to decision outcomes, LLM groups showed diverse profiles, for example, reduced sensitivity to norms or enhanced impartiality. We report model differences in when and how strongly the boost manifests. We also discuss prompt and agent compositions that enhance or mitigate the effect. We end with a discussion of the implications for AI alignment, multi-agent design, and artificial moral reasoning. Code available at: https://github.com/baltaci-r/MoralAgents
@@ 1 Introduction
Multi-agent systems (MAS) provide a key paradigm for decentralized decision-making and coordination among multiple autonomous entities, which is especially helpful in dynamic environments. Incorporating Large Language Models (LLMs), systems referred to as LLM-MAS, feature agents with enhanced capabilities in reasoning, communication, and knowledge access [1]. This integration significantly expands the scope and complexity of tasks MAS can tackle, supporting more adaptive, flexible, and human-aligned agent behaviors [2]. However, integrating LLMs into MAS also introduces new risks. When agents negotiate through iterative message passing or jointly refine solutions, even minor alignment discrepancies can cascade into major emergent distortions. Bias amplification [3], covert coordination protocols [4], knowledge drift [5], conflicting agreements [6], and collusion [7, 8] phenomena may emerge, all of which remain invisible when each LLM agent is evaluated in isolation. As LLM-MAS are deployed to tackle increasingly complex tasks, understanding how collective reasoning emerges from individual models is imperative for ensuring safety and trustworthiness [9].
@@
Figure 1: A schematic representing our experimental setup for LLM moral deliberation and reflection. A triad of LLM agents engages in multi-round discussions about moral dilemmas and concludes with private reflections. This example illustrates how the group setting induces a Utilitarian Boost whereby moral norm violation is endorsed in the service of a "greater good".
@@
Recent work draws on insights from social psychology to investigate emergent distortions in group-level multi-agent LLM reasoning, demonstrating phenomena such as conformity [10, 11], belief congruence [12], and social irrationality [4] (see the Related Works section for a more detailed overview). Although not unique to LLM-MAS, these emergent patterns often show both similarities to and differences from well-studied collective behaviors in humans [13]. Such distortions become especially important in domains where real-world consequences of norm violations can be severe, as in healthcare, education, and law. Aligning individual models will be insufficient to guarantee aligned group outcomes in such cases, and debiasing would need to consider the collective element of decision-making.
Moral reasoning is a type of decision-making increasingly delegated to LLM collectives across high-stakes domains. Examples include coordinating specialized doctor and patient agents in simulated clinical settings [14], and decomposing complex legal tasks to improve legal reasoning and interpretation [15]. However, research on LLM moral reasoning has remained focused on the individual, typically comparing single agents with humans (e.g., see [16, 17, 18]). Therefore, it provides little insight into emergent collective moral dynamics.
@@
This knowledge gap increases the risk of missing important deviations from guidelines, with potentially major real-world consequences. In morally charged settings, emergent social dynamics could lead to collective LLM decisions that contradict those of reasonable humans, any normative frameworks, and even the output of any single agent. This creates an urgent blind spot: without analyzing group-level moral dynamics in LLM-MAS, we cannot understand, predict, or prevent ethically problematic outcomes that escape single-agent safety evaluations. Consider, for instance, a case where clinical LLM groups would endorse sacrificing an individual for "the greater good" in contrast with a physician’s Hippocratic oath. The resulting loss of trust in experts and decision-making institutions would be profound.
Studying group dynamics in moral LLM decision-making also fills an important gap in computational social science, where LLMs are increasingly used for simulating human reasoning or determining whether human tendencies arise from language use [19]. Ignoring the deeply social nature of moral reasoning, in which group deliberation, argumentation, and negotiation play central roles, limits the insights afforded by LLM simulation experiments.
To address these gaps, the current study uses controlled experiments and validated psychological measurement tools to investigate how collective moral reasoning occurs in LLM-MAS. More specifically, it focuses on the following questions:
@@
    Are LLMs more likely to endorse norm violation in the service of a "greater good" when deliberating in groups? We operationalize this "Utilitarian Boost" by placing multiple copies of the same LLM in a structured dialogue, asking them to reach consensus on diverse sets of moral scenarios. The scenarios include difficult dilemmas, for instance, whether a doctor should poison a patient who intends to infect others with a deadly disease.
@@
    How do LLM-MAS judgments compare to human group judgments? Drawing on human data from group deliberation studies, we compare quantitative changes in the endorsement of violations and note qualitative differences between human and LLM responses.
@@
    How does any Utilitarian Boost arise? We compare theories from moral psychology as explanations by examining, for instance, whether sensitivity to norms is reduced within groups, or whether collective settings encourage more impartial judgments of decision outcomes.
  4. 4.
    Can any Utilitarian Boost be amplified or mitigated with LLM-MAS design choices? We evaluate role prompt changes, model pairings, and conversation frames to determine how the boost can be adjusted.
As a preview, we found a concerning phenomenon not previously demonstrated in multi-agent LLM systems: when identical agents engage in collective deliberation, they do not merely aggregate their individual moral judgments. Instead, they exhibit a consistent Utilitarian Boost: a systematic shift toward endorsing actions that maximize overall welfare, even when such actions involve sacrificing or harming a minority of humans. After six rounds of inter-agent exchange, groups reach a utilitarian decision that is more approving of moral norm violations than the corresponding single-agent models. This change appears in six popular LLMs and is therefore unlikely to reflect specific training datasets or paradigms. Reliability checks through human ratings make incoherent responses or misalignment between LLMs’ ratings and their argumentation unlikely explanations. Although a Utilitarian Boost has been observed in human groups [20, 21, 22], we find the LLM-MAS Utilitarian Boost often happens under markedly different circumstances, with significant variation across models.
The Utilitarian Boost in LLM-MAS moral decision-making is particularly concerning given that such systems are already deployed in ethically-sensitive domains, where they offer guidance in emotionally complex situations to resolve dilemmas involving fairness, harm, and moral responsibility [23, 24]. Some researchers even urge their broader deployment by claiming that LLMs can outperform professional ethicists in moral judgment tasks [25]. To help mitigate the risks of such deployments, we discuss a range of tests that examined potential strategies for mitigating the Utilitarian Boost, be it through design choices or diversity in agent roles and models.
@@ 2 Related Work
2.1 Emergent Biases in LLM Multi-Agent Systems
Researchers are increasingly applying insights from social psychology to understand emergent reasoning biases in LLM-MAS. In this vein, Liu et al. [4] introduce CogMir, a framework that leverages systematic hallucinations in LLMs to model human cognitive biases within a collective decision-making context. Their findings reveal both parallels and divergences between human and LLM group behavior, particularly in how agents express prosocial tendencies under uncertainty. Weng et al. [10] investigated conformity among LLM agents, showing that models tend to align their responses with peers under social influence, mirroring human-like conformity effects. Borah et al. [12] examined belief congruence in LLM collectives, where models broadly align themselves with fellow agents who demonstrate shared beliefs. The presence of similarly emergent collective distortions in moral reasoning remains unclear.
@@ 2.2 Moral Reasoning in Human Groups
When humans act as a group, their collective moral reasoning cannot be fully explained by the aggregate attitudes of individual members. When decisions are made collectively, people more readily accept norm breaches that yield greater benefits for a larger number of people [20, 21, 22]. Empirical studies have explored a variety of potential causes for this shift. A heightened sense of rationality in groups could induce a Utilitarian Boost [20]. Alternatively, joint allocation of limited resources in group discussions can lead to greater consideration for the least well-off, making it easier to go against rigid norms [26]. Group deliberation may also reduce the stress and negative emotions associated with norm violations [21]. A sense of social connection among group members can also make it easier to make sacrifices for the "greater good". A recent study used computational modeling to tease apart when and to what extent several of the proposed mechanisms contribute to a Utilitarian Boost in humans [27].They estimate three parameters based on subjects’ responses to moral scenarios: stronger sensitivity to a decision’s consequences (C), weaker sensitivity to norms (N), and preference for passive rather than active violations (I; Inaction preference). Consider the example in Figure 1. A higher value of C would emphasize minimizing harm for all those involved, which would result in considering the poisoning as justified. Lower N would make the moral rule against a physician harming their patient less prominent in the decision. Finally, lower I would reflect a stronger inclination to act rather than remain passive, even when that action entails harm. Estimating these parameters based on responses to moral scenarios, researchers found the Utilitarian Boost in human groups to be driven solely by heightened sensitivity to consequences (C)[22]. These studies may inform the types of moral decision-making distortions we could expect in LLM-MAS. Whether a similar pattern holds for LLM-MAS is unexplored.
@@ 3.1 Experimental Design
As a baseline, we instructed individual LLMs to independently evaluate responses to sets of moral dilemmas as described below (Solo condition). Agents provided moral acceptability ratings where higher scores showed prioritizing the overall welfare of all involved over respect for one’s moral duties (i.e., utilitarianism over deontological reasoning). We then created groups of LLM agents with size $s=\{2,3\}$ who collaboratively engaged with the same dilemmas over discrete rounds of conversation $t=\{1,2,...6\}$, discussing their reasoning and working towards a shared judgment (Group condition). In each round, each agent contributed a detailed response that built on the conversation, reconsidered their position if warranted, and concluded with a moral acceptability rating (henceforth called the utilitarianism score). At the end of the discussion, each agent was asked to privately reflect and generate a detailed, individual argument explaining their final reflection score.
To ensure the validity and alignment of the generated arguments and the accompanying utilitarian scores, we conducted a human evaluation study on Prolific. A stratified sample comprising approximately 1% of the model-generated arguments, balanced by dilemma type, model, and condition, was double-rated on 7-point Likert scales for their degree of utilitarian and deontological support. Each item was rated independently by two crowd-sourced participants who passed an attention check. As detailed in Appendix F, this external validation confirmed that our LLM-generated ratings and justifications align with coherent moral interpretations, ensuring that subsequent group–solo comparisons reflect meaningful moral reasoning rather than superficial linguistic variation.
After confirming the validity of the agent responses, we compared them across conditions and discussion stages using mixed-effects regression models with the <ordinal> package in R[28]. Random intercepts were included to account for variability across dilemmas. We conducted experiments on several moral benchmarks, detailed in Section 3.2, using six popular LLMs: five open-source models (Llama3.3:70B [29], QwQ [30], Qwen3:32B [31], Gemma3:27B [32], and Qwen2.5:32B [33]) and one closed-source model (GPT4.1 [34]). Among these, QwQ and Qwen3 are reasoning models. Each trial was repeated three times to ensure reliability (henceforth called repetitions). Detailed model settings, prompts, and experimental results for all LLMs are given in Appendices A, B, and C, respectively.
@@ 3.2 Stimuli and Measures
To see whether collective deliberation among LLMs alters moral reasoning relative to solo judgments, we use a layered design informed by established methods in moral psychology. We begin with (i) classical sacrificial dilemmas developed by Greene and colleagues, which provide a well-established foundation for studying conflicts between outcome-based (utilitarian) and rule-based (deontological) moral reasoning. These dilemmas have shaped two decades of research on moral cognition and emotion, serving as the tool of choice for identifying Utilitarian Boosts in humans [35, 21]. We complement this set of dilemmas with (ii) measures that reflect different potential causes for the Utilitarian Boost observed in human groups. This multidimensional approach allows us to characterize each LLM’s moral profile, identify which factors contribute to any Utilitarian Boosts within LLM groups, and determine whether they correspond to the causes observed in humans.
However, we are not interested only in demonstrating collective distortions to LLM moral judgments, but also which factors amplify or mitigate them. This information can guide LLM-MAS design and interventions for safer, more trustworthy systems. We therefore conduct (iii) post-hoc probing and mitigation experiments by varying model diversity, replacing multi-agent exchange with structured self-reflection, and prompt-enforced moral reasoning styles to test when and how shifts emerge and dampen.
(i) General Utilitarian Boost.
The most classic tool for evaluating utilitarian moral reasoning in humans consists of many dilemmas that distinguish between personal and impersonal situations, compared with non-moral controls  [36, 35, 37, 38]. Personal dilemmas involve direct, hands-on harm to a person (for instance, by pushing them to their death) that aims to prevent harm to many others. Impersonal dilemmas revolve around more indirectly causing such harm (e.g., by flipping a switch) to the same end. Humans often find it easier to make the utilitarian choice of harming the individual in impersonal scenarios. Large cross-cultural studies show that the Utilitarian Boost for impersonal harm applies widely, supporting the view that this set of dilemmas captures a universal feature of human moral cognition [39]. Prior human group-decision work using the same set of dilemmas shows that collective deliberation increases utilitarian scores relative to making individual judgments, suggesting the scenarios as suitable vehicles to look for similar Utilitarian Boosts in LLMs [21, 20].
Emotional activation is the most favored explanation for the distinct effects of personal and impersonal situations  [35, 40]. Reduced negative emotions when harm is indirect might make it easier for individuals to inflict harm in return for broader benefits  [38, 21]. To see if any Utilitarian Boosts in LLMs were likely learned from emotional language use, we analyzed the emotional tone of model-generated justifications using pretrained emotion classifiers. We applied two complementary models: one trained on six basic emotions plus a neutral category [41], and another trained on Google’s GoEmotions dataset [42], which predicts 27 fine-grained emotional categories. This dual-tagging approach allows us to examine how collective deliberation may alter the emotional framing of a dilemma, therefore affecting the agents’ utilitarianism scores. Details of classifier architectures and preprocessing are described in Appendix E.4.
(ii) Mechanisms of Utilitarian Reasoning.
Beyond the overall utilitarianism score, we probe which components of utilitarian reasoning are affected by collective deliberation using a few complementary instruments. [^1] We initially planned to include the Factual Moral Framework [43] as well, but discovered an implementation error during final validation that prevented reliable results. We report only the validated frameworks here.
Oxford Utilitarianism Scale (OUS). When studying humans, both Impartial Beneficence (equal concern for everyone’s welfare) and Instrumental Harm (endorsement of harm as a means to an end)  [44] encourage violating norms in service of a "greater good". Thus OUS allows us to measure whether a Utilitarian Boost in LLM groups is due to the discussion making their judgments more impartial, increasing tolerance for causing harms in the service of an end, or shifts only one of these components.
CNI Model. The CNI model [27] allows researchers to computationally derive three agent-specific latent variables from scenario responses, distinguishing between different kinds of Utilitarian Boost: An agent’s sensitivity to the Consequences of their decision (C), their level of sensitivity to norms in general (N), and their Inaction preference (I).
Of the three latent variables, N and I can be calculated directly by applying CNI equations to agents’ responses for dilemmas that orthogonally vary norm compliance vs. violation alongside action vs. omission  [27]. This results in four conditions: Action–Congruent, where the moral norm demands taking an action, Action–Incongruent, where norm dictates not performing an action, Omission–Congruent, where norm suggests allowing something to happen, and Omission–Incongruent, where norm forbids allowing an outcome through inaction. If a surgeon terminates a comatose patient to use their organs for saving five others, we are faced with an Action-Incongruent situation, as medical norms forbid actively harming a patient to whatever end.
The third latent variable, C, is calculated from the combination of the answers to the four sets of scenarios. More specifically, if agents more often choose action when its benefits exceed the costs, controlling for norm sensitivity and general action bias, sensitivity to consequences will be higher. Conversely, if their choices vary little with the cost-benefit manipulation, the value of C will be lower.
(iii) Post-hoc Probing and Mitigation. After establishing the existence of Utilitarian Boosts across models in LLM-MAS, we conducted a series of post-hoc experiments to see how they can be parametrically modulated. We systematically varied three dimensions:
  1. 1.
    Agent and model diversity. We paired different models and divergent LLM parameters, testing whether model "diversity" alters the group’s Utilitarian Boost.
  2. 2.
    Self-reflection depth. It is possible that longer moral reflection causes a Utilitarian Boost, regardless of whether a group deliberation is involved. We tested this possibility by performing an additional experiment where multi-agent discussions were replaced with self-debates, in which a single model iteratively critiqued and revised its own reasoning.
  3. 3.
    Prior seeding. We enforced divergent moral reasoning styles (deontological, utilitarian, and neutral) in LLM agents paired together to test how prior moral framing shapes convergence.
These manipulations allow us to probe the mechanisms underlying the group shift and identify potential levers to mitigate undesirable Utilitarian Boosts, while keeping the dilemma content and scoring identical to those in the main experiment.